The ratio of 1.2 to 1 is equal to which of the following ratios?
(A) 1 to 2
(B) 12 to 1
(C) 5 to 6
(D) 6 to 5
(E) 6 to 50

Answer by stanbon(75887)

The ratio of 1.2 to 1 is equal to which of the following ratios?
(A) 1 to 2
(B) 12 to 1
(C) 5 to 6
(D) 6 to 5
(E) 6 to 50
----------
1.2/1 = 12/10 = 6/5
6 to 5
=================
Cheers,
Stan H.